# Beyond Words: On Large Language Models Actionability in Mission-Critical Risk Analysis

### v2

@@ 1. Introduction
We base our findings on the experience and data acquired during late autumn 2023 and early spring 2024. We leveraged FTM and RAG-assisted models. Notably, HEs demonstrated higher accuracy, while LLMs are quicker and more actionable. Moreover, our findings show that RAG-assisted LLMs have the lowest hallucination rates, effectively uncovering hidden risks and complementing human expertise. Thus, the choice of model depends on specific needs, with FTMs for accuracy and RAG models for comprehensiveness and actionability. Therefore, experts can focus on more comprehensive risk analysis, leveraging LLMs for a practical complementing companion in risk analysis within a condensed timeframe. They can also save costs by averting unnecessary expenses associated with implementing unwarranted countermeasures.
Our study’s main contributions are as follows:
@@ 2. Background and Definitions
In this section, we provide essential concepts to our study, encompassing the RAG and fine-tuning procedures, as well as the definitions of vulnerability, threat, scenario, and sample within our specific context.
@@ 2.1. Definitions
We define “scenario” as an excerpt of an RA interview, likewise, a “sample” includes the scenario along with the description of the associated risk, if any.
According to National Security Authority (1995), we define threat (in Italian “minaccia”) as “the possibility of accidental or deliberate compromise of the security of a computer system or network, resulting in the loss of information confidentiality, data modification/destruction, or interruption or alteration of service behavior(National Security Authority, 1995).
@@ 2.2. Retrieve-Augmented Generation
The knowledge of LLM models is of two types: Parametric and Non-Parametric. It learns parametric knowledge during training, implicitly storing it in the neural network’s weights. Conversely, non-parametric knowledge resides in external sources such as a vector database. A vector database stores data as high-dimensional vectors, representing features or attributes mathematically. An example of Non-parametric knowledge is the one provided by RAG (Lewis et al., 2020) that allows the enhancement of the capabilities of LLMs by using the user question to the model as a query for its vector database and sending back to the model as contextual information, the result of the search and the original user question.
The vector database store information as embeddings (Lewis et al., 2020). Figure 1 presents the embedding workflow. Documents are analyzed through a specialized model (e.g. OpenAi’s text-embedding-3-large  [^1] https://platform.openai.com/docs/guides/embeddings/embedding-models), and the resulting word embedding are stored in the vector store.
@@ 2.4. Risk Analysis Process
The RA process requires acquiring information regarding the current state of the art about safety and security measures implemented to guard sensitive information, how we select and train the personnel involved, and the procedures adopted in their treatment.
Such information can be internal documents, best practices, technical blueprints, and interviews.
The personnel must be part of the internal Information Security Department, and a certain number of workers, depending on the dimension of the company or public administration, must be randomly chosen from all the people who treat sensitive information.
Then, all the gathered material is analyzed and confronted with what the applicable law provides regarding possible threats that insist on sensitive information given the scenario, the corresponding vulnerabilities, and the suitable countermeasures. All of this is to calculate a threshold index that we must at least match to have reasonable confidence that the implemented measures can assure the security of our sensitive information.
@@ 3. Related Works
Spiess et al. (2024) emphasize the significance of calibration in LLM, especially for tasks like code generation in SE. They introduce a framework for evaluating the calibration of code-generating models and find that these models are typically not well-calibrated initially. The authors propose methodologies, like Platt scaling, to enhance calibration, aiming to improve decision-making regarding the use of model-generated code.
Ni et al. (2023) developed ChatReport, an LLM-based system, to automate the analysis of corporate sustainability reports. They recognized the challenge of limited resources for human analysis due to the dense and complex nature of these reports, leading to a lack of transparency in sustainability reporting. To address this, they aimed to democratize sustainability report analysis by empowering stakeholders with automatic analysis tools. However, they encountered challenges such as LLM hallucination and the inefficiency of involving domain experts in AI development. Our results aligns with previous studies (Martino et al., 2023).
@@
Esposito et al. (Esposito and Palagiano, 2024) investigated the proficiency of LLM in PSRA. Our work deepens Esposito et al.’s (Esposito and Palagiano, 2024) investigation extending the analysis to the more comprehensive RA. Therefore, we investigate the comprehensiveness and actionability of the LLM beyond simple accuracy.
@@ 4.3. Data Analysis
To answer $RQ_{1}$, we measured the accuracy of the four models on the 357 samples provided.
We adopted standard accuracy metrics in IR as suggested by (Esposito and Palagiano, 2024; Chang et al., 2023).
We tested $H_{01}$ to assess whether using RAG improves the LLM accuracy and proficiency by comparing their outputs against our GT.
To answer $RQ_{2}$, we measured the actionability and comprehensiveness of the model’s RA.
Although GT comes from finalized RAs, we considered the possibility for human experts to overlook some details, thus failing to identify specific or more subtle threats. We call this threat “hidden risks”. Therefore, we also evaluated the ability of models and human experts to extend the ground truth (GT) to measure if the finalized RAs (GT) were spotless or if hidden risks were present. Table 3 presents the metrics we used to evaluate the scenarios. More specifically, we instructed the HRs to analyze the HE and the model’s output according to Table LABEL:tab:prediction_types and Table 3.
@@ 5. Results
This Section presents our study results and findings to answer our RQs.
@@ 5.4. Comparing LLMs with Human Experts (RQ3)
Table 11 shows WT p-values. According to Table 11 we can reject the null hypothesis for the metrics Extended GT and Actionable. Therefore, we can affirm that the output of the LLMs model is more actionable than HEs analyses.
@@
Table 11. P-value and gain of the scenario metrics differences, i.e. H04 test results.
Metrics P-Value Gain
Actionable 0,0031 65%
Comprehensive 0,2312 12%
Extended GT 0,0115 28%
@@
Finally, considering our findings, we can affirm that LLMs can outperform humans in actionability, hidden risk discovery and analysis speed. Nonetheless, LLMs fall short in accuracy. Hence, future works must focus on improving LLMs’ RA accuracy.
@@ 6.2. Model Explainability
In our study, we tasked HR to measure the actionability of the LLM’s output. According to Table 10, LLMs provided consistently actionable outputs.
Therefore, according to Arrieta et al. (2020), we define our audience as the RA experts, including those who analyze the risk and assess the report. As a metric to measure explainability, we use the actionability of the model.
@@ 7. Industrial Implications
In this section, we discuss industrial-specific implications.
The pervasive integration of technology across various domains of human activity, coupled with escalating aggression from attackers driven by fraudulent, demonstrative, or political motives, has resulted in a surge not only in the development of intrusion tools and techniques but also in an exponential rise in the number and diversity of attacks. (Clusit, [n. d.])
The dynamic nature of the scenario clashes with the lengthy execution times of processes like risk analysis, which entail complex steps such as data collection, analysis, evaluation, countermeasure identification, resource allocation, procedural adjustments, and staff training.
@@
The idea of applying LLMs to speed-up evaluation phases by up to 92% significantly presents a compelling opportunity. This efficiency improvement, coupled with the model’s superior performance achieving a 65% enhancement in actionability, a 12% increase in comprehensiveness, and a 28% extension of the GT compared to HEs, raises two key implications:
@@
    we can develop a training process for junior evaluators that incorporates plausible scenarios created by senior experts. This process would assist junior evaluators in comparing their answers with those generated by the model, enabling them to identify and understand their errors. By doing so, they can enhance their skills ”in-house” without solely relying on extensive fieldwork, although fieldwork remains valuable and should not be completely eliminated.
@@ 8. Threats to Validity
Construct Validity concerns how our measurements reflect what we claim to measure (Wohlin et al., 2012). Our specific design choices may impact our results, including our measurement process and data filtering. To address this threat, we based our choice on past studies and used well-established guidelines in designing our methodology (Runeson and Höst, 2009; Basili et al., 1994).

### v4

@@ 1. Introduction
We base our findings on the experience and data acquired during late autumn 2023 and early spring 2024. We leveraged FTM and RAG-assisted models. Notably, HEs demonstrated higher accuracy, while LLMs are quicker and more actionable. Moreover, our findings show that RAG-assisted LLMs have the lowest hallucination rates, effectively uncovering hidden risks and complementing human expertise. Thus, the choice of model depends on specific needs, with FTMs for accuracy and RAG models for comprehensiveness and actionability. Therefore, experts can focus on more comprehensive risk analysis, leveraging LLMs for a practical complementing companion in risk analysis within a condensed timeframe. They can also save costs by averting unnecessary expenses associated with implementing unwarranted countermeasures. Our study’s main contributions are as follows:
@@ 2. Background and Definitions
In this section, we provide essential concepts to our study, encompassing the RAG and fine-tuning procedures and the definitions of vulnerability, threat, scenario, and sample within our specific context.
@@ 2.1. Definitions
We define “scenario” as an excerpt of an RA interview. Likewise, a “sample” includes the scenario along with the description of the associated risk, if any. According to National Security Authority (1995), we define threat (in Italian “minaccia”) as “the possibility of accidental or deliberate compromise of the security of a computer system or network, resulting in the loss of information confidentiality, data modification/destruction, or interruption or alteration of service behavior(National Security Authority, 1995).
@@ 2.2. Retrieve-Augmented Generation
The knowledge of LLM models is of two types: Parametric and Non-Parametric. It learns parametric knowledge during training, implicitly storing it in the neural network’s weights. Conversely, non-parametric knowledge resides in external sources such as a vector database. A vector database stores data as high-dimensional vectors, representing features or attributes mathematically. An example of Non-parametric knowledge is the one provided by RAG (Lewis et al., 2020) that allows the enhancement of the capabilities of LLMs by using the user question to the model as a query for its vector database and sending back to the model as contextual information, the result of the search and the original user question. The vector database store information as embeddings (Lewis et al., 2020). Figure 1 presents the embedding workflow. Documents are analyzed through a specialized model (e.g. OpenAi’s text-embedding-3-large  [^1] https://platform.openai.com/docs/guides/embeddings/embedding-models), and the resulting word embedding are stored in the vector store.
@@ 2.4. Risk Analysis Process
Figure 3 presents the RA workflow. RA process (National Security Authority, 1995) requires acquiring information (Phase 1) regarding the current state of the art about safety and security measures implemented to guard sensitive information, how we select and train the personnel involved, and the procedures adopted in their treatment (Step 1). Such information can be internal documents, best practices, technical blueprints, and interviews.
The personnel must be part of the internal Information Security Department, and a certain number of workers, depending on the dimension of the company or public administration, must be randomly chosen from all the people who treat sensitive information (National Security Authority, 1995; for Standardization, 2022).
We then assess the data gathered to identify the threats to CIA (Confidentiality, Integrity, Availability)(of Standards and Technology, 2020) (Step 2) that can be exploited through the vulnerabilities found (Step 3). We assign an initial vulnerability value (Step 4) given the scenario.
In the next step (Phase 2), all the gathered material is analyzed and presented with what the applicable laws, best practices, and technology provide regarding possible countermeasures to the identified threats and vulnerabilities (Step 5).
Following this, we compute the residual vulnerability value, which results after applying all the countermeasures that can be implemented in the system (Step 6).
All of this is to calculate a threshold index (named “P”) that we must at least match to have reasonable confidence that the implemented measures can assure the security of our sensitive information (Step 7).
If the “P” index is equal to or less than 30, we must identify more countermeasures from Step 5. At the same time, if it’s above 30, we can proceed with the last part of the RA process (Phase 3), where we discuss all the findings of the possible countermeasures and compute the percentage weight of technical countermeasures compared to the total number of countermeasures identified (Step 8) (the higher the better (National Security Authority, 1995)) to find the equivalent level of information assurance (Step 9), both in regard of ITSEC(National Security Authority, 1995; European Commission and Media, 1991) and EAL(Criteria, 2022) standards.
@@ 3. Related Works
Ni et al. (2023) developed ChatReport, an LLM-based system, to automate the analysis of corporate sustainability reports. They recognized the challenge of limited resources for human analysis due to the dense and complex nature of these reports, leading to a lack of transparency in sustainability reporting. However, they encountered challenges such as LLM hallucination and the inefficiency of involving domain experts in AI development. Our results aligns with previous studies (Martino et al., 2023).
@@
Finally, Esposito et al. (Esposito and Palagiano, 2024) investigated the proficiency of LLM in PSRA. Our work deepens Esposito et al.’s (Esposito and Palagiano, 2024) investigation, extending the analysis to the more comprehensive RA.Referring to Figure 3, the difference between PSRA (Esposito and Palagiano, 2024) and RA (our work) is that the first one is limited to the first 3 steps, while RA covers the entire process. Therefore, we investigate the comprehensiveness and actionability of the LLM beyond simple accuracy.
@@ 4.3. Data Analysis
To answer $RQ_{1}$, we measured the accuracy of the four models on the 357 samples provided. We adopted standard accuracy metrics in IR as suggested by (Esposito and Palagiano, 2024; Chang et al., 2023). We tested $H_{01}$ to assess whether using RAG improves the LLM accuracy and proficiency by comparing their outputs against our GT.
To answer $RQ_{2}$, we measured the actionability and comprehensiveness of the model’s RA. Although GT comes from finalized RAs, we considered the possibility for human experts to overlook some details, thus failing to identify specific or more subtle threats. We call this threat “hidden risks”. Therefore, we also evaluated the ability of models and human experts to extend the ground truth (GT) to measure if the finalized RAs (GT) were spotless or if hidden risks were present. Table 3 presents the metrics we used to evaluate the scenarios. More specifically, we instructed the HRs to analyze the HE and the model’s output according to Table LABEL:tab:prediction_types and Table 3.
@@ 5. Results
This Section presents our study results and answers our RQs.
@@ 5.4. Comparing LLMs with Human Experts (RQ3)
Regarding $H_{04}$, we can reject the null hypothesis for the metrics Extended GT (p-value 0.0115, gain 28%) and Actionable (p-value 0.0031, gain 65%). Therefore, we can affirm that the output of the LLMs model is more actionable than HEs analyses. While we cannot reject it for the Comprehensive metrics (p-value 0.2312, gain 12%).
@@
Finally, considering our findings, we can affirm that LLMs can outperform humans in actionability, hidden risk discovery, and analysis speed. Nonetheless, LLMs fall short in accuracy. Hence, future works must focus on improving LLMs’ RA accuracy.
@@ 6.2. Model Explainability
In our study, we tasked HR to measure the actionability of the LLM’s output. According to Table 10, LLMs provided consistently actionable outputs. Therefore, according to Arrieta et al. (2020), we define our audience as the RA experts, including those who analyze the risk and assess the report. As a metric to measure explainability, we use the actionability of the model.
@@ 7. Industrial Implications
In this section, we discuss industrial-specific implications. The pervasive integration of technology across various domains of human activity, coupled with escalating aggression from attackers driven by fraudulent, demonstrative, or political motives, has resulted in a surge not only in the development of intrusion tools and techniques but also in an exponential rise in the number and diversity of attacks. (Clusit, [n. d.]) The dynamic nature of the scenario clashes with the lengthy execution times of processes like risk analysis, which entail complex steps such as data collection, analysis, evaluation, countermeasure identification, resource allocation, procedural adjustments, and staff training.
@@
Referring to Figure 3, aside from the initial step of identifying the resources to protect, our approach, leveraging LLMs, can significantly aid and expedite the subsequent steps of the RA workflow. Applying LLMs to speed up evaluation phases by up to 92% presents a compelling opportunity. This efficiency improvement, coupled with the model’s superior performance achieving a 65% enhancement in actionability, a 12% increase in comprehensiveness, and a 28% extension of the GT compared to HEs, raises two key implications:
@@
    we can develop a training process for junior evaluators incorporating plausible scenarios created by senior experts. This process would assist junior evaluators in comparing their answers with those generated by the model, enabling them to identify and understand their errors. By doing so, they can enhance their skills ”in-house” without solely relying on extensive fieldwork, although fieldwork remains valuable and should not be eliminated.
@@ 8. Threats to Validity
Construct Validity concerns how our measurements reflect what we claim to measure (Wohlin et al., 2012). Our design choices, including our measurement process and data filtering, may impact our results. To address this threat, we based our choice on past studies and well-established guidelines in designing our methodology (Runeson and Höst, 2009; Basili et al., 1994).